sum of three consecutive numbers is 18.what are these numbers?
Answer by user_dude2008(1825)

"sum of three consecutive numbers is 18." ----> x+(x+1)+(x+2)=18

x+(x+1)+(x+2)=18

x+x+1+x+2=18

3x+3=18

3x=15

x=5

--> x+1=6, x+2=7

Ans: nums are 5, 6, and 7

Check: 5+6+7=18